OpenStudy (anonymous):

This is my last question. Please help :) Thanks!<333 (medals are given) What is the radius of a circle with a circumference of 23.55 meters? Use 3.14 for pi and show your work.

OpenStudy (dumbcow):

C = 2pi*r solve for radius --> r = C/2pi

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 2\pi r = 23.55 }\) \(\Large \color{MidnightBlue}{\Rightarrow 2 * 3.14 * r = 23.55 }\) Do it now?

OpenStudy (anonymous):

Circumference of a circle = 2 x Pi x Radius :. Radius = Circumference / 2 x Pi = 23.55 / (2 x 3.14) = 3.75

OpenStudy (anonymous):

OMG THANK YOU!

OpenStudy (anonymous):

np :D